A 225 gram of mercury has a volume of 16.5 ml. What is the density of mercury? What is its specific gravity?



1
Expert's answer
2021-08-02T12:27:12-0400

Solution:

Density of the object = Mass / Volume

Therefore,

Density of mercury = (225 g) / (16.5 mL) = 13.636 g/mL = 13.6 g/mL

Density of mercury = 13.6 g/mL


Specific gravity = Density of the object / Density of water

Density of water = 1 g/mL

Therefore,

Specific gravity = (13.6 g/mL) / (1 g/mL) = 13.6

Specific gravity = 13.6


Answer:

The density of mercury is 13.6 g/mL

Specific gravity is 13.6